OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)?

OpenStudy (mathmale):

To find f(x) * g(x), simply indicate multiplication of log x and (5x+2):\[f(x)*g(x) = (\log x)*(5x+2)\] That's it. No need to carry out the actual multiplication (unless your teacher asks you to do so).

OpenStudy (anonymous):

Yeah they actually want me to multiply it out everything i come out with is wrong these are the answer choices f(x) • g(x) = log10 x5x − 2 f(x) • g(x) = log10 (5x − 2)x f(x) • g(x) = 5x log10 x + 2 log10 x f(x) • g(x) = 2 log10 x − 5x log10 x